The 2017 balance sheet of Dream, Inc., showed current assets of $4,500 and current liabilities of $2,920. The 2018 balance sheet showed current assets of $3,000 and current liabilities of $1,530.

What was the company’s 2018 change in net working capital, or NWC? (Do not round intermediate calculations. A negative answer should be indicated by a minus sign.)
Answer #1

Net Working Capital = Current Assets - Current Liabilities

Net Working Capital in 2017 = 4,500-2,920 = $1,580

Net working capital in 2018 = $3,000-$1,530 = $1,470

company’s 2018 change in net working capital, or NWC = 1,470-1,580

= -$110